Nitrous acid in 0.1 F solution is 5.4 percent ionized. Calculate the ionization constant for this acid.


1
Expert's answer
2021-03-12T06:02:18-0500

Nitrous acid dissociation:

HNO2 + H2O = H3O+ + NO2-

Ionization (dissociation) constant can be calculated using the formula:

Ka = [H3O+] [NO2-] / [HNO2]

Since the solution is 5.4 % ionized:

[H3O+] = [NO2-] = 5.4 % × 0.1 F / 100% = 0.0054 F

[HNO2] = 0.1 F - 0.0054 F = 0.0946 F

From here:

Ka = [H3O+] [NO2-] / [HNO2] = 0.0054 × 0.0054 / 0.0946 = 3.08 × 10-4


Answer: 3.08 × 10-4